OpenStudy (twizttiez):

What is the approximate volume of a sphere with a radius of 9 in.? Use 3.14 to approximate pi and express your answer in hundredths. ??? in^3

OpenStudy (twizttiez):

@clynnew

OpenStudy (sleepyjess):

Do you know the formula for volume of a sphere?

OpenStudy (twizttiez):

Can't say I do :/

OpenStudy (sleepyjess):

\(V = \dfrac 43 \pi r^3\)

OpenStudy (twizttiez):

Would It be 305.63??

OpenStudy (twizttiez):

@sleepyjess

OpenStudy (sleepyjess):

sorry, was afk for a minute I am getting a lot bigger

OpenStudy (twizttiez):

Bigger? XD

OpenStudy (sleepyjess):

yah

OpenStudy (twizttiez):

Like how so?

OpenStudy (sleepyjess):

way way bigger

OpenStudy (twizttiez):

Wait like are you getting tagged a lot?

OpenStudy (sleepyjess):

no haha, the answer I'm getting is a lot bigger than what you got

OpenStudy (twizttiez):

Oh XD

OpenStudy (twizttiez):

Well I rounded it to the nearest hundredth

OpenStudy (sleepyjess):

how did you get 305.63???

OpenStudy (twizttiez):

The google volume calculator :/

OpenStudy (twizttiez):

I'm not gonna lie XD

OpenStudy (sleepyjess):

Okay, first, what's our radius?

OpenStudy (twizttiez):

9

OpenStudy (sleepyjess):

Okay, so now we have \(\dfrac 43 \pi 9^3\) We are supposed to use 3.14 for \(\pi\) What is 4/3?

OpenStudy (twizttiez):

I got 3052.08

OpenStudy (sleepyjess):

Yep :)

OpenStudy (twizttiez):

Oh but wouldn't I round it to 305.08?

OpenStudy (sleepyjess):

That is 1/10th of the actual number...

OpenStudy (twizttiez):

Then what do I round it to??

OpenStudy (sleepyjess):

3 0 5 2 . 0 8 thousands hundreds tens ones tenths hundreths

OpenStudy (twizttiez):

Ohhhhhhhh :| sorry

OpenStudy (sleepyjess):

It's already rounded since we only went out to the hundreths place in pi :)

OpenStudy (twizttiez):

Ohhhhhh okay so then could you recheck like 4 more of my questions??

OpenStudy (sleepyjess):

sure :)

OpenStudy (twizttiez):

Thanks so much

OpenStudy (sleepyjess):

I'm going to put my work here for a question really quick if that's okay with you

OpenStudy (twizttiez):

Yeah no prob